how do you solve the question y=2x-3
Answer by Mathtut(3670)

you can make a table and graph it...choose any number for x or y and solve for the remaining variable. such as
if x =2 then y=2(2)-3=1 so you have the ordered pair(2,1) as one point
if x=3 then y=2(3)-3=3..............................(3,3) as another point....with 2 points you can graph this. Plot more points to verify you line is correct.